Use differentials to approximate the volume of material needed to make a rubber ball if the radius of the hollow inner core is 2 in., and the thickness of the rubber is 1/8 in.


1
Expert's answer
2021-11-23T09:44:02-0500

"t=2 \space in"

"\sigma_c=500 Mpa"

"\tau=200 Mpa"

Solution:

as we know that 

"\sigma_c\times A_c = \tau\times A_s"

"\sigma_c\times\pi/4 D^2=\tau\times\pi D t"

"500/4 (D)= 200\times 10"

"D= 16 mm"